# An Analytical and Experimental Study of Distributed Uplink Beamforming in the Presence of Carrier Frequency Offsets

### v1

@@ Abstract
Index Terms:
MU-MIMO, distributed multi-user beamforming (D-MUBF), carrier frequency offset (CFO), estimation error
@@ I Introduction
Distributed multiple-input multiple-output (MIMO) systems [1, 2, 3, 4, 5] have long been recognized for their potential to enhance coverage range, spectral efficiency, and energy efficiency of wireless networks. Distributed MIMO takes advantage of multiple multi-antenna access points (APs) across a shared coverage area, allowing them to jointly serve a large number of users. For a network of distributed APs (dAPs) to efficiently serve multiple users on the same time-frequency resource, synchronization between the dAPs is essential [3, 6, 4, 7, 8]. In infrastructure networks, utilizing high-speed, error-free optical fiber to connect dAPs can offer precise synchronization by sharing a common clock. However, this approach may not always be practical due to the varying network topologies and the high costs of deployment. Likewise, in ad hoc networks, such as unmanned aerial vehicles (UAVs) or sensor networks, a reliable and error-free link between distributed nodes is often not feasible. As a result, over-the-air (OTA) synchronization remains a crucial step for enabling distributed multi-user MIMO (MU-MIMO) transmissions. This underscores the importance of evaluating system performance across different network topologies when employing OTA synchronization methods.
@@
OTA synchronization faces several challenges that significantly impact system performance [9, 10, 11]. In [9], Ganesan et al. addressed the challenges of OTA carrier synchronization in distributed multi-antenna systems, emphasizing the impact of the accuracy of synchronization algorithms on beamforming performance. In general, the accuracy of OTA synchronization methods is inherently dependent on channel conditions and the accuracy of local oscillators (LOs). A summary of recent developments in carrier frequency offset (CFO) estimation techniques can be found in [12]. A low signal-to-noise ratio (SNR) regime or bad channel conditions reduces the accuracy of CFO estimation, leading to inter-user interference and reduced spectral efficiency. Therefore, residual frequency offsets, resulting either from imperfections in the synchronization process or LO drifts, are of significant concern. The problem is further aggravated in distributed scenarios where the data transmitted or received by dAPs is affected by the combined effect of frequency offsets between all users and dAPs, making the traditional CFO compensation methods less effective. Thus, it is critical to characterize the impact of the uncompensated CFO among dAPs on the performance of distributed multi-user beamforming (D-MUBF).
@@
Beamforming has been extensively studied in the context of MU-MIMO systems and distributed antenna systems, particularly with a focus on achieving optimal spectral efficiency and enhancing wireless network capacity [13, 14, 15, 16]. Most of these works do not consider non-ideal synchronization and imperfect CFO estimation and lack comprehensive experimental validations. Early studies on MU-MIMO beamforming in centralized architectures assumed ideal synchronization among APs and users [15]. However, achieving accurate synchronization in distributed systems has proven to be more challenging due to the inherent independence of LOs and the distributed nature of the antennas [3]. To enable distributed MU-MIMO transmission in time division duplex (TDD) mode, the practical design and implementation of synchronization techniques have been investigated for various network topologies, including wireless local area network (WLAN) [3], distributed MIMO systems [17, 2, 1, 18], and UAV networks [19]. More recently, receiver designs have been proposed to jointly estimate CFO, channel, and data, in a MU-MIMO setting based on the orthogonal frequency division multiplexing (OFDM) by assuming Gaussian-mixture distributions for phase offsets caused by CFO [20]. However, these approaches are typically evaluated through simulations or numerical analysis under idealized conditions, and they often remain susceptible to practical impairments, such as residual CFO due to LO drifts and estimation errors, particularly in low-SNR regimes. These imperfections can severely degrade beamforming performance in real-world deployments.
@@
The analytical characterization of hardware impairments and synchronization errors—particularly CFO—and their impact on system performance has been extensively studied across various wireless systems, including WLAN [3], massive MIMO [21, 22, 23], and millimeter wave (mmWave) systems [24]. For instance, a closed-form expression for SNR degradation in the presence of CFO for OFDM-based index modulation systems is derived in [25]. Although valuable, these studies do not address the impact of the statistical properties of CFO in distributed systems on beamforming. Moreover, they lack the extensive experimental validation needed to fully understand the real-world implications of residual CFO in distributed systems. Experimental measurements and field trials have been widely performed in several studies, including 5G-based distributed MU-MIMO [26, 27], 5G-based distributed MIMO in the 28 GHz band [28], distributed beamforming with drones [29], and distributed beamforming with indoor massive MIMO base station [30], among others. Most of these empirical measurements are based on a specific protocol stack or network topology and do not consider the random residual CFO in distributed systems. To the best of our knowledge, there is a lack of experimentally validated analytical studies that specifically examine the impact of residual inter-dAP CFOs on the performance of D-MUBF.
@@
In this paper, we fill this gap by introducing an analytical framework that studies the beamforming performance as a function of CFO statistics, thus offering a more comprehensive understanding of D-MUBF performance in the presence of random residual CFOs. We also validate our framework with experimental measurements performed using the RENEW massive MIMO testbed [31] with perfect and imperfect synchronization. In particular, our paper outlines three main contributions, as detailed in the following.
@@
Second, we conduct extensive experiments using the RENEW massive MIMO testbed [31] in an indoor environment on the Rice University campus to evaluate D-MUBF techniques in the presence of CFO. We collected over 250 datasets, each containing up to 2000 frames, across various experimental scenarios. These scenarios include MU-MIMO configurations emulating both distributed and co-located AP deployments, with and without perfect synchronization. In the distributed scenario, each dAP in the testbed operates with an independent LO to derive carrier and sampling frequencies, resulting in random inter-dAP CFOs. Within each dataset, every frame includes both channel sounding pilots and multi-user uplink data intended for receive beamforming. The datasets are publicly available online [32], and we offer insights on utilizing them for future research endeavors.
@@ II System Model with Carrier Frequency Offsets
We consider a network of $M$ APs with $N$ antennas each, deployed in a given coverage area serving $K$ single-antenna users. It is assumed that APs are connected to a central processing unit, which can be viewed as edge cloud processors [33]. To illustrate the reference network architecture, Fig. 1 presents an example consisting of two regions, each with a cluster of dAPs connected to a local processor and further linked to a central server (CS), jointly serving multiple users. The CS controls cluster formation, scheduling, precoding, and beamforming operations. The dAPs in each cluster work collaboratively to serve all the users in the coverage area. This paper considers a fixed cluster of $M$ dAPs simultaneously serving $K$ users, without addressing the problem of cluster formation. The transmission process follows the TDD standard, where uplink pilot, uplink data, and downlink data transmissions are separated in the time domain.
In a cluster of $M$ dAPs, for each transmission attempt, the CS determines one of the dAPs as the Master dAP, and schedules $K$ users to be served on the same time-frequency resources. The master dAP is responsible for broadcasting a synchronization beacon to the $K$ scheduled users. The users will then use this beacon to synchronize their frame start time with the dAPs and schedule their uplink pilot. After transmitting orthogonal uplink pilots, all $K$ users start transmitting uplink data simultaneously. An uplink slot comprising $N_{\mathrm{slot}}^{\mathrm{ul}}$ OFDM symbols is considered for data transmission. Following that, a downlink slot can be considered during which dAPs transmit precoded downlink data to the $K$ users. For downlink transmission, a reciprocity calibration step is necessary to compensate for the non-reciprocal effects of the hardware [34]. In this work, we only consider uplink transmission to remove the need for hardware calibration and focus exclusively on the impact of CFO on beamforming performance.
@@ II-A Carrier Frequency Offset Model
In a MU-MIMO setup, multiple users share the uplink slot for simultaneous data transmission. The received signal is affected by both the user-to-dAP channel conditions and the combined impact of synchronization imperfections across each user-dAP pair. Synchronization between a user transmitter and a dAP receiver ensures alignment in both frame timing and carrier frequency, enabling OFDM demodulation with negligible ICI. This can be achieved using well-established synchronization schemes widely adopted in current cellular and WLAN systems [3, 35]. Conversely, in distributed MU-MIMO systems, synchronization across dAPs is critical, as each dAP operates with an independent LO with unique CFO characteristics. Since the received signals from all cooperating dAPs are combined through receive beamforming to demultiplex user signals, precise inter-dAP synchronization is required to compensate for the relative phase rotations among dAPs.
We assume that although dAPs perform an OTA synchronization procedure, a residual relative CFO remains at each dAP. This residual offset stems from imperfections in the OTA synchronization process—affected by channel conditions—and from LO drifts, which are influenced by hardware variability, temperature fluctuations, aging, and other environmental factors. The residual frequency offset between dAPs disrupts subcarrier orthogonality, leading to ICI, and introduces phase shifts across successive OFDM symbols. These phase shifts accumulate during receive combining, ultimately degrading the resulting receive SINR.
@@ II-B OFDM in the Presence of CFO
$\textstyle\mathrm{\check{y}}_{m}^{(i)}[n,\zeta]=\frac{1}{\sqrt{N_{sc}}}\,e^{j2\pi(\zeta+n(N_{sc}+L_{cp}))T_{s}\delta_{f,m}}\,$
@@
in which $\zeta$ indicates the index of the time-domain OFDM samples, $q$ represents the subcarrier index, and $T_{s}$ is the symbol time. The frequency-domain representation of the received symbol is obtained by applying the FFT to the time-domain signal in (II-B), with the full derivation provided in Appendix A. Accordingly, the frequency-domain (noise-free) received signal is given by
@@
$\overline{\phi}_{m}[n]=2\pi\left(1+\frac{L_{cp}}{N_{sc}}\right)n\epsilon_{m}=2\pi\left(N_{sc}+L_{cp}\right)nT_{s}\delta_{f,m},$ (6)
@@ III-A Uplink Multi-User Beamforming
It is considered in this analysis that all $N$ antennas at all $M$ dAPs are participating in beamforming. Previous simulations presented in [30] indicated that in distributed uplink beamforming, where there is no perfect frequency synchronization between dAPs, increasing the number of receive antennas in beamforming does not necessarily lead to higher received SNR or reduced error vector magnitude (EVM). Notably, the experimental results in [30], evaluating receive beamforming using zero-forcing and conjugate methods (Figs. 4 and 5, respectively), demonstrate that in the presence of CFO, increasing the number of beamforming antennas from 8 to 64 provides no improvement in EVM for the single-user scenario and only limited improvement for multi-user scenarios.
Using the estimated signals in (III-A), the average uplink SINR for a given user $k$ can be derived, when all $M$ dAPs are receiving the uplink signals and cooperating in beamforming. By considering the channel vectors from different users uncorrelated, e.g., due to enough spatial separation and rich scattering, and by defining the transmit SNR of each user as $\gamma_{k}=p_{k}/\sigma^{2}_{z}$, the average uplink SINR can be expressed as
@@
In D-MUBF methods, beamforming weights are derived based on channel estimates from users to mitigate inter-user interference [14]. The performance of the D-MUBF methods depends on several factors, including number of users, number of antennas at each dAP, spatial correlation between antennas, and residual phase and frequency offsets [13]. D-MUBF can be implemented using various beamforming techniques. In this work, we focus on two widely used approaches: conjugate (CBF) [36] and zero-forcing beamforming (ZFBF) [37]. These methods are selected due to their practical relevance and complementary characteristics—CBF is favored for its simplicity and robustness to channel estimation errors, while ZFBF provides higher interference suppression at the cost of increased computational complexity. Evaluating both allows us to assess D-MUBF performance across a representative trade-off space between implementation complexity and interference mitigation. To better understand the impact of CFO on SINR, we derive a conditional closed-form expression for the SINR, evaluated for fixed channel vectors and expressed as a function of the statistical properties of the inter-dAP CFO component. This enables the evaluation of SINR with respect to the inter-dAP CFO distribution for any given channel realization.
@@ IV-A Numerical Evaluation of the SINR Expression
The SINR expression in (III-B2) is a function of channel vectors and the expectations of the CFO component in the forms of $\mathbb{E}\{\Omega\}$ and $\mathbb{E}\{|\Omega|^{2}\}$. We evaluate the asymptotic behavior of the SINR with respect to the standard deviation (STD) of the normalized CFO component, defined in (3). The expectation term $\mathbb{E}\{\Omega\}$ for both CFO distributions—uniform (given in (19)) and normal (given in (21))—decays and converges to zero as the STD increases (i.e., $\alpha/\sqrt{3}$ for uniform and $\beta$ for normal), and also with increasing OFDM symbol index $n$ in the time domain. In addition, from (20) and (22), it can be inferred that $\left(\mathbb{E}\{\Omega\}\right)^{2}$ approaches zero when $n\alpha/\sqrt{3}>0.25$ for the uniform distribution and $n\beta>0.25$ for the normal distribution. Therefore, within these specified ranges of $n$, $\alpha$, and $\beta$, the term $\left(\mathbb{E}\{\Omega\}\right)^{2}$ can be asymptotically excluded from the SINR expression in (III-B2). We note that the distribution parameters of the normalized CFO term in (3) are chosen to ensure the validity of the approximation introduced in (16).

### v2

@@ I Introduction
Distributed multiple-input multiple-output (MIMO) systems [airsync, megamimo, rogalin, rashid_2022_syncAlg, Rashid_2023_syncAlg] have long been recognized for their potential to enhance coverage range, spectral efficiency, and energy efficiency of wireless networks. Distributed MIMO takes advantage of multiple multi-antenna access points (APs) across a shared coverage area, allowing them to jointly serve a large number of users. For a network of distributed APs (dAPs) to efficiently serve multiple users on the same time-frequency resource, synchronization between the dAPs is essential [rogalin, martirosov_2022_syncAlg, rashid_2022_syncAlg, ahmad_2024_syncAlg, milos_2023_syncAlg]. In infrastructure networks, utilizing high-speed, error-free optical fiber to connect dAPs can offer precise synchronization by sharing a common clock. However, this approach may not always be practical due to the varying network topologies and the high costs of deployment. Likewise, in ad hoc networks, such as unmanned aerial vehicles (UAVs) or sensor networks, a reliable and error-free link between distributed nodes is often not feasible. As a result, over-the-air (OTA) synchronization remains a crucial step for enabling distributed multi-user MIMO (MU-MIMO) transmissions. This underscores the importance of evaluating system performance across different network topologies when employing OTA synchronization methods.
@@
OTA synchronization faces several challenges that significantly impact system performance [beamsync, yao_sync_stat, palat_ber_upper_bound]. In [beamsync], Ganesan et al. addressed the challenges of OTA carrier synchronization in distributed multi-antenna systems, emphasizing the impact of the accuracy of synchronization algorithms on beamforming performance. In general, the accuracy of OTA synchronization methods is inherently dependent on channel conditions and the accuracy of local oscillators (LOs). A summary of recent developments in carrier frequency offset (CFO) estimation techniques can be found in [10702556]. A low signal-to-noise ratio (SNR) regime or bad channel conditions reduces the accuracy of CFO estimation, leading to inter-user interference and reduced spectral efficiency. Therefore, residual frequency offsets, resulting either from imperfections in the synchronization process or LO drifts, are of significant concern. The problem is further aggravated in distributed scenarios where the data transmitted or received by dAPs is affected by the combined effect of frequency offsets between all users and dAPs, making the traditional CFO compensation methods less effective. Thus, it is critical to characterize the impact of the uncompensated CFO among dAPs on the performance of distributed multi-user beamforming (D-MUBF).
@@
Beamforming has been extensively studied in the context of MU-MIMO systems and distributed antenna systems, particularly with a focus on achieving optimal spectral efficiency and enhancing wireless network capacity [aryanfar_mu_bf, zafari2024asilomar, argos2012, keerthi_ul_mu_mimo]. Most of these works do not consider non-ideal synchronization and imperfect CFO estimation and lack comprehensive experimental validations. Early studies on MU-MIMO beamforming in centralized architectures assumed ideal synchronization among APs and users [argos2012]. However, achieving accurate synchronization in distributed systems has proven to be more challenging due to the inherent independence of LOs and the distributed nature of the antennas [rogalin]. To enable distributed MU-MIMO transmission in time division duplex (TDD) mode, the practical design and implementation of synchronization techniques have been investigated for various network topologies, including wireless local area network (WLAN) [rogalin], distributed MIMO systems [chorus, megamimo, airsync, airshare], and UAV networks [airbeam]. More recently, receiver designs have been proposed to jointly estimate CFO, channel, and data, in a MU-MIMO setting based on the orthogonal frequency division multiplexing (OFDM) by assuming Gaussian-mixture distributions for phase offsets caused by CFO [10806522]. However, these approaches are typically evaluated through simulations or numerical analysis under idealized conditions, and they often remain susceptible to practical impairments, such as residual CFO due to LO drifts and estimation errors, particularly in low-SNR regimes. These imperfections can severely degrade beamforming performance in real-world deployments.
@@
The analytical characterization of hardware impairments and synchronization errors, particularly CFO, and their impact on system performance has been extensively studied across various wireless systems, including WLAN [rogalin], massive MIMO [Emil_Non_Ideal_HW2015, zhang2018freq_sync, ulf2014hw_impairments], and millimeter wave (mmWave) systems [chatelier_phase_noise_mmwave]. For instance, a closed-form expression for SNR degradation in the presence of CFO for OFDM-based index modulation systems is derived in [besseghier2024performance]. Although valuable, these studies do not address the impact of the statistical properties of CFO in distributed systems on beamforming. Moreover, they lack the extensive experimental validation needed to fully understand the real-world implications of residual CFO in distributed systems. Experimental measurements and field trials have been widely performed in several studies, including 5G-based distributed MU-MIMO [theoni2020dist_mu_mimo5g, shinya2019exp_trial5g], 5G-based distributed MIMO in the 28 GHz band [daisuke2017outdoor_exp5g], distributed beamforming with drones [junming2019exp_demo], and distributed beamforming with indoor massive MIMO base station [sam2022], among others. Most of these empirical measurements are based on a specific protocol stack or network topology and do not consider the random residual CFO in distributed systems. To the best of our knowledge, there is a lack of experimentally validated analytical studies that specifically examine the impact of residual inter-dAP CFOs on the performance of D-MUBF.
@@
In this paper, we fill this gap by introducing an analytical framework that studies the beamforming performance as a function of CFO statistics, thus offering a more comprehensive understanding of D-MUBF performance in the presence of random residual CFOs. We also validate our framework with experimental measurements performed using the RENEW massive MIMO testbed [renew] with perfect and imperfect synchronization. In particular, our paper outlines three main contributions, as detailed in the following.
@@
Second, we conduct extensive experiments using the RENEW massive MIMO testbed [renew] in an indoor environment on the Rice University campus to evaluate D-MUBF techniques in the presence of CFO. We collected over 250 datasets, each containing up to 2000 frames, across various experimental scenarios. These scenarios include MU-MIMO configurations emulating both distributed and co-located AP deployments, with and without perfect synchronization. In the distributed scenario, each dAP in the testbed operates with an independent LO to derive carrier and sampling frequencies, resulting in random inter-dAP CFOs. Within each dataset, every frame includes both channel sounding pilots and multi-user uplink data intended for receive beamforming. The datasets are publicly available online [dist_mu_mimo_datasets], and we offer insights on utilizing them for future research endeavors.
@@ II System Model with Carrier Frequency Offsets
We consider a network of $M$ APs with $N$ antennas each, deployed in a given coverage area serving $K$ single-antenna users. It is assumed that APs are connected to a central processing unit, which can be viewed as edge cloud processors [c-ran]. To illustrate the reference network architecture, Fig. 1 presents an example consisting of two regions, each with a cluster of dAPs connected to a local processor and further linked to a central server (CS), jointly serving multiple users. The CS controls cluster formation, scheduling, precoding, and beamforming operations. The dAPs in each cluster work collaboratively to serve all the users in the coverage area. This paper considers a fixed cluster of $M$ dAPs simultaneously serving $K$ users, without addressing the problem of cluster formation. The transmission process follows the TDD standard, where uplink pilot, uplink data, and downlink data transmissions are separated in the time domain.
In a cluster of $M$ dAPs, for each transmission attempt, the CS determines one of the dAPs as the Master dAP, and schedules $K$ users to be served on the same time-frequency resources. The master dAP is responsible for broadcasting a synchronization beacon to the $K$ scheduled users. The users will then use this beacon to synchronize their frame start time with the dAPs and schedule their uplink pilot. After transmitting orthogonal uplink pilots, all $K$ users start transmitting uplink data simultaneously. An uplink slot comprising $N_{\mathrm{slot}}^{\mathrm{ul}}$ OFDM symbols is considered for data transmission. Following that, a downlink slot can be considered during which dAPs transmit precoded downlink data to the $K$ users. For downlink transmission, a reciprocity calibration step is necessary to compensate for the non-reciprocal effects of the hardware [shepard2020argosnet]. In this work, we only consider uplink transmission to remove the need for hardware calibration and focus exclusively on the impact of CFO on beamforming performance.
@@ II-A Carrier Frequency Offset Model
In a MU-MIMO setup, multiple users share the uplink slot for simultaneous data transmission. The received signal is affected by both the user-to-dAP channel conditions and the combined impact of synchronization imperfections across each user-dAP pair. Synchronization between a user transmitter and a dAP receiver ensures alignment in both frame timing and carrier frequency, enabling OFDM demodulation with negligible ICI. This can be achieved using well-established synchronization schemes widely adopted in current cellular and WLAN systems [rogalin, schmidl1997]. Conversely, in distributed MU-MIMO systems, synchronization across dAPs is critical, as each dAP operates with an independent LO with unique CFO characteristics. Since the received signals from all cooperating dAPs are combined through receive beamforming to demultiplex user signals, precise inter-dAP synchronization is required to compensate for the relative phase rotations among dAPs.
We assume that although dAPs perform an OTA synchronization procedure, a residual relative CFO remains at each dAP. This residual offset stems from imperfections in the OTA synchronization process, affected by channel conditions, and from LO drifts, which are influenced by hardware variability, temperature fluctuations, aging, and other environmental factors. The residual frequency offset between dAPs disrupts subcarrier orthogonality, leading to ICI, and introduces phase shifts across successive OFDM symbols. These phase shifts accumulate during receive combining, ultimately degrading the resulting receive SINR.
@@ II-B OFDM in the Presence of CFO
$\textstyle\mathrm{\check{y}}_{m}^{(i)}[n,\zeta]=\frac{1}{\sqrt{N_{sc}}}\,e^{j\frac{2\pi}{N_{sc}}(\zeta+n(N_{sc}+L_{cp}))\epsilon_{m}}$
@@
in which $\zeta$ indicates the index of the time-domain OFDM samples and $q$ represents the frequency-domain subcarrier index. The frequency-domain representation of the received symbol is obtained by applying the FFT to the time-domain signal in (II-B), with the full derivation provided in Appendix A. Accordingly, the discrete-time noise-free received signal in the frequency-domain is given by
@@
$\overline{\phi}_{m}[n]=\frac{2\pi}{N_{sc}}\left(N_{sc}+L_{cp}\right)n\epsilon_{m}=2\pi\left(1+\frac{L_{cp}}{N_{sc}}\right)n\epsilon_{m},$ (6)
@@ III-A Uplink Multi-User Beamforming
It is considered in this analysis that all $N$ antennas at all $M$ dAPs are participating in beamforming. Previous simulations presented in [sam2022] indicated that in distributed uplink beamforming, where there is no perfect frequency synchronization between dAPs, increasing the number of receive antennas in beamforming does not necessarily lead to higher received SNR or reduced error vector magnitude (EVM). Notably, the experimental results in [sam2022], evaluating receive beamforming using zero-forcing and conjugate methods (Figs. 4 and 5, respectively), demonstrate that in the presence of CFO, increasing the number of beamforming antennas from 8 to 64 provides no improvement in EVM for the single-user scenario and only limited improvement for multi-user scenarios.
Using the estimated signals in (III-A), the average uplink SINR for a given user $k$ can be derived when all $M$ dAPs are receiving the uplink signals and cooperating in beamforming. By considering the channel vectors from different users uncorrelated, e.g., due to enough spatial separation and rich scattering, and by defining the transmit SNR of each user as $\gamma_{k}=p_{k}/\sigma^{2}_{z}$, the average uplink SINR can be expressed as
@@
In D-MUBF methods, beamforming weights are derived based on channel estimates from users to mitigate inter-user interference [zafari2024asilomar]. The performance of the D-MUBF methods depends on several factors, including number of users, number of antennas at each dAP, spatial correlation between antennas, and residual phase and frequency offsets [aryanfar_mu_bf]. D-MUBF can be implemented using various beamforming techniques. In this work, we focus on two widely used approaches: conjugate (CBF) [giovanni2021cbf] and zero-forcing beamforming (ZFBF) [hong2025zfbf]. These methods are selected due to their practical relevance and complementary characteristics—CBF is favored for its simplicity and robustness to channel estimation errors, while ZFBF provides higher interference suppression at the cost of increased computational complexity. Evaluating both allows us to assess D-MUBF performance across a representative trade-off space between implementation complexity and interference mitigation. To better understand the impact of CFO on SINR, we derive a conditional closed-form expression for the SINR, evaluated for fixed channel vectors and expressed as a function of the statistical properties of the inter-dAP CFO component. This enables the evaluation of SINR with respect to the inter-dAP CFO distribution for any given channel realization.
@@ IV-A Numerical Evaluation of the SINR Expression
The SINR expression in (III-B2) is a function of channel vectors and the expectations of the CFO component in the forms of $\mathbb{E}\{\Omega\}$ and $\mathbb{E}\{|\Omega|^{2}\}$. We evaluate the asymptotic behavior of the SINR with respect to the standard deviation (STD) of the normalized CFO component, defined in (3). The expectation term $\mathbb{E}\{\Omega\}$ for both CFO distributions, uniform (given in (19)) and normal (given in (21)), decays and converges to zero as the STD increases (i.e., $\alpha/\sqrt{3}$ for uniform and $\beta$ for normal), and also with increasing OFDM symbol index $n$ in the time domain. In addition, from (20) and (22), it can be inferred that $\left(\mathbb{E}\{\Omega\}\right)^{2}$ approaches zero when $n\alpha/\sqrt{3}>0.25$ for the uniform distribution and $n\beta>0.25$ for the normal distribution. Therefore, within these specified ranges of $n$, $\alpha$, and $\beta$, the term $\left(\mathbb{E}\{\Omega\}\right)^{2}$ can be asymptotically excluded from the SINR expression in (III-B2). We note that the distribution parameters of the normalized CFO term in (3) are chosen to ensure the validity of the approximation introduced in (16).